# OSUM: Advancing Open Speech Understanding Models with Limited Resources in Academia

### v1

@@ Introduction
In this study, we aim to foster broader academic exploration of SULMs with limited resource demands, encouraging wider research community participation. To this end, we introduce OSUM, an open SULM with its data processing pipeline and training details publicly available. The OSUM model integrates a Whisper speech encoder, fine-tuned on a multi-task dataset, with a Qwen2 LLM. It is capable of performing a wide range of speech tasks, including automatic speech recognition (ASR), speech recognition with timestamps (SRWT), vocal event detection (VED), speech emotion recognition (SER), speaking style recognition (SSR), speaker gender classification (SGC), speaker age prediction (SAP), and speech-to-text chat (STTC). Notably, SSR is a distinctive feature of our OSUM model and serves as a vital component of speech understanding. It enhances the model’s capability by improving contextual comprehension and boosting performance across various downstream speech tasks. Furthermore, it establishes a foundation for enabling more natural and context-aware speech-based interactions. We adopt an ASR+X training strategy to enhance training stability and reduce resource consumption for our SLUM model, wherein an auxiliary ASR task is optimized alongside the primary target task (denoted as “X”). For instance, during the training of the SER task, we concurrently train the ASR task (ASR+SER) by predicting both transcription and emotion labels for each speech sample. This multi-task training accelerates modality alignment, enabling the LLM to effectively utilize both textual and acoustic modalities. Our OSUM model utilizes only 44,100 hours of training data and achieves comparable or superior performance to other SULMs. The overall performance of OSUM is illustrated in Fig. 1. The model is trained on Nvidia A6000 GPUs and Huawei Ascend NPUs, supporting inference on both platforms. The goal of this study is to foster transparency and accelerate progress in the field of SULMs by providing accessible tools and resources for the broader research community.
@@ 2.3 Training Data
Table 2: Details of the multi-task training data for OSUM. The total duration is 43,100 hours. The total training data includes both open-sourced datasets and the data we processed.
@@ 2.4 Data Processing Pipe-line
Efforts to train a speaker gender classification model to label web-sourced data yield unsatisfactory performance. Consequently, we discard the pseudo-labeled data and relied solely exclusively on manually labeled datasets for training. For the SGC task, we select KeSpeech (Tang et al., 2021), Datatang-Kid (Datatang Tech Co Ltd, 2024), AISHELL-1 (Bu et al., 2017), AISHELL-2 (Du et al., 2018), LibriSpeech (Panayotov et al., 2015), Kaggle-CommonVoice (Kaggle Community, 2017), and Magicdata-Read OpenSLR (2019) as training dataset, as they include reliable speaker gender labels.
@@
Similar to the SGC task, due to the poor performance of the automated labeling model, only manually labeled data is used for training. We use KeSpeech (Tang et al., 2021), Datatang-Kid (Datatang Tech Co Ltd, 2024), Magicdata-Read (OpenSLR, 2019), Kaggle-CommonVoice (Kaggle Community, 2017), AISHELL-ASR0060 (AISHELL Tech Co Ltd, 2024), and AISHELL-ASR0018 (AISHELL Tech Co Ltd, 2024) as the training dataset for the SAP task, as these datasets provide reliable speaker age labels.
@@ 3.1 Training Setup
In the second stage, we conduct experiments on 24 Huawei Ascend NPUs, using a learning rate of 5e-5. The process completes a total of 528,000 training steps and consumes 7.5 days.
@@ 3.3 Main Results
As illustrated in Table 4, our approach reveals obvious advantages in the ASR task on the Chinese test sets. Notably, the proposed OSUM consistently outperforms other models on the WenetSpeech test-meeting set, three AISHELL-2 sub-test sets, and three internally used SpeechIO test sets. While OSUM does not surpass the top-performing method on the English test set, it rivals performance comparable to SenseVoice-S. These results are achieved, remarkably, with substantially less training data. This underscores that our ASR+X task paradigm effectively enhances model convergence in ASR tasks, significantly minimizing the data and computational resources required for SULMs training.
@@
Table 5 presents the SRWT evaluation results for our proposed OSUM model compared to Whisper-Large-v3, Qwen-Audio, and the GMM-HMM model used for generating annotated data in SRWT tasks. Our OSUM model significantly outperforms Whisper-Large-v3 by relative ${31.38\%}$ and also surpasses Qwen-Audio. Additionally, our OSUM’s SRWT performance is very close to the GMM-HMM model, which is widely regarded for its high accuracy in timestamp prediction. These results underscore the effectiveness of OSUM in the SRWT task. While OSUM’s performance is slightly below that of the GMM-HMM model, its primary advantage lies in its ability to predict timestamps in an end-to-end fashion, simplifying the integration with other tasks, such as speaker diarization.
@@
In the SGC task, we evaluate Qwen2-Audio and OSUM. The results demonstrate that OSUM achieves an ${100\%}$ accuracy on the AISHELL-1 test set. While this result is commendable, we suspect it may indicate some degree of overfitting. Furthermore, on the Kaggle test set, our approach slightly outperforms Qwen2-Audio, yet it falls short on our internal test set. This is likely due to the presence of background music in our internal test set, a characteristic less prevalent in OSUM’s training data. Addressing this limitation is a focus of our future work. Overall, OSUM exhibits its value in the SGC task.
@@
We also compare our OSUM with Qwen2-Audio on the SAP task. During our previous experiments, we found that the acoustic similarity between teenagers and adults is remarkably high, complicating effective differentiation. Consequently, we categorize age into three groups: child, adult, and old. Curiously, despite our efforts to debug the prompts, Qwen2-Audio demonstrates a lower age classification accuracy on both the Kaggle test set and our internal test set. This may stem from their overly detailed age categorization, which hinders the model’s training accuracy. Our model significantly surpasses Qwen2-Audio on the Kaggle test set, achieving an accuracy of ${76.72\%}$. Although the classification accuracy slightly declines on our proprietary test set, it still outperforms Qwen2-Audio. This indicates that our model exhibits strong generalization capabilities on different data.
@@
In the STTC task, we follow AirBench’s evaluation protocol across all test sets. This involves providing the text of audio queries along with the text of two distinct answers, allowing a text-based LLM to assign subjective scores from 1 to 10. The two answers consist of a real response and the answer generated by SULMs. While AirBench employs GPT-4 as the scoring LLM, it is currently inaccessible, so we instead utilize GPT-3.5-Turbo. The test results presented in Table 5 indicate that, on AirBench’s official speech sub-test set, our score is lower than that of Qwen2-Audio, suggesting that our model’s capabilities in English conversation and audio description lag behind those of Qwen2-Audio. This is primarily because we did not use English conversational data for training; the current score relies entirely on the LLM’s performance. Furthermore, the gap narrows significantly on our internal Chinese conversational test set, suggesting that our strategy of performing ASR before chat is beneficial. Overall, our OSUM still lags behind Qwen2-Audio in conversational ability, which will be a key focus of our future improvements.

### v2

@@ Introduction
In this study, we aim to foster broader academic exploration of SULMs with limited resource demands, encouraging wider research community participation. To this end, we introduce OSUM, an open SULM with its data processing pipeline and training details publicly available. The OSUM model integrates a Whisper speech encoder, fine-tuned on a multi-task dataset, with a Qwen2 LLM. It is capable of performing a wide range of speech tasks, including automatic speech recognition (ASR), speech recognition with timestamps (SRWT), vocal event detection (VED), speech emotion recognition (SER), speaking style recognition (SSR), speaker gender classification (SGC), speaker age prediction (SAP), and speech-to-text chat (STTC). Notably, SSR is a distinctive feature of our OSUM model and serves as a vital component of speech understanding. It enhances the model’s capability by improving contextual comprehension and boosting performance across various downstream speech tasks. Furthermore, it establishes a foundation for enabling more natural and context-aware speech-based interactions. We adopt an ASR+X training strategy to enhance training stability and reduce resource consumption for our SLUM model, wherein an auxiliary ASR task is optimized alongside the primary target task (denoted as “X”). For instance, during the training of the SER task, we concurrently train the ASR task (ASR+SER) by predicting both transcription and emotion labels for each speech sample. This multi-task training accelerates modality alignment, enabling the LLM to effectively utilize both textual and acoustic modalities. Our OSUM model utilizes only 50,500 hours of training data and achieves comparable or superior performance to other SULMs. The overall performance of OSUM is illustrated in Fig. 1. The model is trained on Nvidia A6000 GPUs and Huawei Ascend NPUs, supporting inference on both platforms. The goal of this study is to foster transparency and accelerate progress in the field of SULMs by providing accessible tools and resources for the broader research community.
@@ 2.3 Training Data
Table 2: Details of the multi-task training data for OSUM. The total duration is 50,500 hours. The total training data includes both open-sourced datasets and the data we processed.
@@ 2.4 Data Processing Pipe-line
Efforts to train a speaker gender classification model to label web-sourced data yield unsatisfactory performance. Consequently, we discard the pseudo-labeled data and relied solely exclusively on manually labeled datasets for training. For the SGC task, we select KeSpeech (Tang et al., 2021), Datatang-Kid (Datatang Tech Co Ltd, 2024), AISHELL-1 (Bu et al., 2017), AISHELL-2 (Du et al., 2018), LibriSpeech (Panayotov et al., 2015), Kaggle-CommonVoice (Kaggle Community, 2017), and Magicdata-Read OpenSLR (2019) as training dataset, as they include reliable speaker gender labels. In addition, we employ a noisy data augmentation strategy.
@@
Similar to the SGC task, due to the poor performance of the automated labeling model, only manually labeled data is used for training. We use KeSpeech (Tang et al., 2021), Datatang-Kid (Datatang Tech Co Ltd, 2024), Magicdata-Read (OpenSLR, 2019), Kaggle-CommonVoice (Kaggle Community, 2017), AISHELL-ASR0060 (AISHELL Tech Co Ltd, 2024), and AISHELL-ASR0018 (AISHELL Tech Co Ltd, 2024) as the training dataset for the SAP task, as these datasets provide reliable speaker age labels. Noisy data augmentation is also utilized in this task.
@@ 3.1 Training Setup
In the second stage, we conduct experiments on 24 Huawei Ascend NPUs, using a learning rate of 5e-5. The process completes a total of 704,000 training steps and consumes 10 days.
@@ 3.3 Main Results
As illustrated in Table 4, our approach reveals obvious advantages in the ASR task on the Chinese test sets. Notably, the proposed OSUM consistently outperforms other models on the WenetSpeech test-meeting set, three AISHELL-2 sub-test sets, and four internally used SpeechIO test sets. While OSUM does not surpass the top-performing method on the English test set, it rivals performance comparable to SenseVoice-S. These results are achieved, remarkably, with substantially less training data. In addition, we find that OSUM exhibits a surprisingly impressive ability to recognize Chinese-English code-mixed speech, even though such code-mixed datasets are not included during training. To be specific, the MER/CER/WER is ${3.89\%}$/${2.41\%}$/${19.30\%}$ on the ASRU code-switching test set (Shi et al., 2020). Going forward, we will enhance this function. Overall, these results underscore that our ASR+X task paradigm effectively enhances model convergence in ASR tasks, significantly minimizing the data and computational resources required for SULMs training.
@@
Table 5 presents the SRWT evaluation results for our proposed OSUM model compared to Whisper-Large-v3, Qwen-Audio, and the GMM-HMM model used for generating annotated data in SRWT tasks. Our OSUM model significantly outperforms Whisper-Large-v3 by relative ${36.70\%}$ and also surpasses Qwen-Audio. Additionally, our OSUM’s performance in the SRWT task even slightly surpasses that of the GMM-HMM model, which is widely recognized for its high accuracy in timestamp prediction. These results underscore the effectiveness of OSUM in the SRWT task. Additionally, OSUM’s high performance in the SRWT task not only enables it to predict timestamps in an end-to-end manner but, more importantly, simplifies its integration with other tasks, such as speaker diarization.
@@
In the SGC task, we evaluate Qwen2-Audio and OSUM. The results demonstrate that OSUM achieves an ${100\%}$ accuracy on the AISHELL-1 test set. While this result is commendable, we suspect it may indicate some degree of overfitting. Furthermore, on the Kaggle test set, our approach slightly outperforms Qwen2-Audio, yet it falls short on our internal test set. This indicates that Qwen2-Audio has strong robustness in the SGC task, which may be due to the fact that they have used a wider range of data. In the follow-up, we plan to continue to increase the training data for this task. Overall, OSUM exhibits its value in the SGC task.
@@
We also compare our OSUM with Qwen2-Audio on the SAP task. During our previous experiments, we found that the acoustic similarity between teenagers and adults is remarkably high, complicating effective differentiation. Consequently, we categorize age into three groups: child, adult, and old. Curiously, despite our efforts to debug the prompts, Qwen2-Audio demonstrates a lower age classification accuracy on both the Kaggle test set and our internal test set. This may stem from their overly detailed age categorization, which hinders the model’s training accuracy. Our model significantly surpasses Qwen2-Audio on the Kaggle test set, achieving an accuracy of ${76.52\%}$. Although the classification accuracy slightly declines on our proprietary test set, it still outperforms Qwen2-Audio. This indicates that our model exhibits strong generalization capabilities on different data.
@@
In the STTC task, we follow AirBench’s evaluation protocol across all test sets. This involves providing the text of audio queries along with the text of two distinct answers, allowing a text-based LLM to assign subjective scores from 1 to 10. The two answers consist of a real response and the answer generated by SULMs. While AirBench employs GPT-4 as the scoring LLM, it is currently inaccessible, so we instead utilize GPT-3.5-Turbo. The test results presented in Table 5 indicate that, on AirBench’s official speech sub-test set, our score is lower than that of Qwen2-Audio, suggesting that our model’s capabilities in English conversation and audio description lag behind those of Qwen2-Audio. This is primarily because we do not use English conversational data for training; the current score relies entirely on the LLM’s performance. However, on our internal Chinese conversational test set, OSUM outperforms Qwen2-Audio, which indicates that our strategy of performing ASR before chat is beneficial. Overall, our OSUM model is comparable to Qwen2-Audio in terms of conversational ability. We will not be content with this achievement. Our future work will mainly focus on the conversation task.